Grade 12th passAlgebra

The real part of each root of the equation (z+1)^6 + z^6=0 is

Prerona Sabui
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
 

Any solution to the equation zn = (z+1)n satisfies |z| = |z+1| i.e. the point is equidistant from (0,0) and (-1,0).

 

This means it lies on the perpendicular bisector of the line joining (0,0) and (-1,0) i.e. x = -1/2 or 2x+1 = 0.

 

Regards

Arun